# Adaptive Width Neural Networks

### v4

@@ Appendix F Dataset Info and Statistics
The synthetic datasets DoubleMoon, Spiral, and SpiralHard are shown in Figure 8. These binary classification datasets have been specifically created to analyze the behavior of AWNN in a controlled scenario where we are sure that the difficulty of the task increases. We provide the code to generate these datasets in the supplementary material.
Figure 8: The DoubleMoon, Spiral, and SpiralHard synthetic datasets are used to test AWNN’s inner workings, and are ordered difficulty. Orange and blue colors denote different classes.
Table 2 reports information on their base statistics and the data splits created to carry out model selection (inner split) and risk assessment (outer split), respectively.
@@
DoubleMoon 2500 2 2 1750/250/500 1800/200
Spiral 2500 2 2 1750/250/500 1800/200
SpiralHard 5000 2 2 3500/500/1000 3600/400
MNIST 70000 28x28 10 53000/7000/10000 54000/6000
CIFAR10 60000 32x32x3 10 44000/6000/10000 45000/5000
CIFAR100 60000 32x32x3 100 44000/6000/10000 45000/5000
NCI1 4110 37 2 3288/411/411 3329/370
REDDIT-B 2000 1 2 1800/100/100 1710/190
Multi30k 31000 128x50257 Multilabel 29000/1000/1000 29000/1000
Note that for Multi30k, the maximum sequence length is 128 and the number of possible tokens is 50527. The classification loss is computed by comparing the predicted token against the expected one, ignoring any token beyond the length of the actual target. For a detailed description of the graph datasets, the reader can refer to Errica et al. [16].
Appendix G Hyper-parameters for fixed and AWNN models
For each data domain, we discuss the set of hyper-parameters tried during model selection for the fixed baseline and for AWNN. We try our best to keep architectural choices identical with the exception of the width of the MLPs used by each model. In all experiments, we either run patience-based early stopping [46] or simply select the epoch with best validation score.
@@ G.1 Synthetic Tabular Datasets and Permuted MNIST
We report the hyper-parameter configurations tried for the fixed MLP/RNN and AWNN in Table 3. In particular, the starting width of AWNN under the quantile function evaluated at $k=0.9$ is approximately $256$ hidden neurons. We also noticed that a learning rate value of $0.1$ makes learning the width unstable, so we did not use it in our experiments. The total number of configurations for the fixed models is 180, for AWNN is 18. We also did not impose any prior on an expected size to see if AWNN recovers a similar width compared to the best fixed model (please refer to Table 1) starting from the largest configuration among the fixed models.
Table 3: Hyper-parameter configurations for standard MLP/RNN and AWNN versions on tabular datasets.
Hyper-Parameter DoubleMoon Spiral SpiralHard
Batch Size 32 $[32,128]$ (MLP/RNN), 128 (AWNN) $[32,128]$ (MLP/RNN), 128 (AWNN)
@@
Optimizer Adam, learning rate $\in[0.1,0.01]$ (MLP/RNN), $0.01$ (AWNN)
AWNN Specific
@@ G.2 Image Classification Datasets
Table 4: Hyper-parameter configurations for standard MLP and AWNN versions on image classification datasets.
@@
In this case, we test 5 different configurations of the fixed baseline, whereas we do not have to perform any model selection for AWNN. The rate of the exponential distribution has been chosen to have a starting width of approximately $128$ neurons, which is the median value among the ones tried for the fixed model.
@@ G.3 Text Translation Tasks
Due to the cost of training a Transformer architecture from scratch on a medium size dataset like Multi30k, we use most hyper-parameters’ configurations of the base Transformer in Vaswani et al. [58], with the difference that AWNN is applied to the MLPs in each encoder and decoder layer. We train an architecture with 6 encoder and 6 decoder layers for 500 epochs, with a patience of 250 epochs applied to the validation loss. We use an Adam optimizer with learning rate 0.01, weight decay 5e-4, and epsilon 5e-9. We also introduce a scheduler that reduces the learning rate by a factor 0.9 when a plateau in the loss is reached. The number of attention heads is set to 8 and the dropout to 0.1. We try different widths for the MLPs, namely [128,256,512,1024,2048]. The embedding size is fixed to 512 and the batch size is 128. The AWNN version starts with an exponential distribution rate of 0.004, corresponding to approximately 512 neurons, and a $\sigma^{\alpha}_{\ell}$=1.
@@ G.4 Graph Classification Tasks
We train the AWNN version of Graph Isomorphism Network, following the setup of Errica et al. [16] and reusing the published results. We test 12 configurations as shown in the table below. Please note that an extra MLP with 1 hidden layer is always placed before the sequence of graph convolutional layers.
Table 5: Hyper-parameter configurations for AWNN versions on graph classification datasets.
@@ Appendix H Re-training a fixed network with the learned width
While on DoubleMoon AWNN learns a similar number of neurons chosen by the model selection procedure and the convergence to the perfect solution over 10 runs seem much more stable, this is not the case on Spiral, SpiralHard and REDDIT-B, where AWNN learns a higher total width over layers. To investigate whether the width learned by AWNN on these datasets is indeed the reason behind the better performances (either a more stable training or better accuracy), we run again the experiments on the baseline networks, this time by fixing the width of each layer to the average width over hidden layers learned by AWNN and reported in Table 1. We call this experiment “Fixed+”.
@@
The results suggest that by using more neurons than those selected by grid search – possibly because the validation results were identical between different configurations – it is actually possible to improve both accuracy and stability of results on Spiral and REDDIT-B, while we had a convergence issue on SpiralHard. Notice that the number of neurons learned by AWNN on REDDIT-B was well outside of the range investigated by Xu et al. [65] and Errica et al. [16], demonstrating that the practitioner’s bias can play an important role on the final performance.
@@ Appendix I Further qualitative results on image classification datasets
Because it is well known that MNIST, CIFAR10, and CIFAR100 are datasets of increasing difficulty, we attach a figure depicting the learned number of neurons and the convergence of AWNN on these datasets, following the practices highlighted in He et al. [26] and described in Table 4 to bring the ResNet20 to convergence. Akin to the tabular datasets, we see a clear pattern where the learned number of neurons, averaged over the 10 final training runs, increases together with the task difficulty. The AWNN version of the ResNet20 converges in the same amount of epochs as the fixed one.
@@
Appendix J Non-linearity ablation study
@@
Appendix K Learned width for each layer on Graph Datasets
@@
Appendix L Comparison with Neural Architecture Search (NAS)
Although complementary to our approach, we thought instructive to run a some basic NAS methods on the image classification datasets. Most NAS methods fix an upper bound on the maximum width, whereas AWNN’s methodology does not. The truncation ability of AWNN is not intended as a substitute for pruning but rather as a confirmation of our claims. In fact, NAS and pruning can actually be combined with AWNN and do not necessarily be seen as “competing” strategies.
The results are shown in Table 6. That said, we compared the performance of AWNN against simple NAS methods on the vision datasets. We tested grid search, random search (Following Wu et al. [64]), local search [60], and Bayesian Optimization [1], fixing the same budget of 5 for all methods and width values between 0 (Linear classifier) and 512 neurons. Results are comparable to those of AWNN.
Table 6: Comparison between different NAS approaches and AWNN. We remind the reader that NAS is an orthogonal and complementary research direction to AWNN.
Grid Search Random Search Local Search Bayes. Optim. AWNN
MNIST 99.6(0.1) 99.6(0.0) 99.5 99.4 99.7(0.0)
CIFAR10 91.4(0.2) 90.1(0.5) 90.6 91.2 91.4(0.2)
CIFAR100 66.5(0.4) 64.9(1.1) 64.9 65.9 63.1(4.0)

### v5

@@ Appendix F Dataset Info and Statistics
The synthetic datasets DoubleMoon, Spiral, and SpiralHard are shown in Figure 8. These binary classification datasets have been specifically created to analyze the behavior of AWN in a controlled scenario where we are sure that the difficulty of the task increases. We provide the code to generate these datasets in the supplementary material.
Figure 8: The DoubleMoon, Spiral, and SpiralHard synthetic datasets are used to test AWN’s inner workings, and are ordered difficulty. Orange and blue colors denote different classes.
Table 3 reports information on all datasets’ statistics and the data splits created to carry out model selection (inner split) and risk assessment (outer split), respectively.
@@
DoubleMoon 5000 2 2 3600/400/1000 3600/400
Spiral 5000 2 2 3600/400/1000 3600/400
SpiralHard 10000 2 2 7200/800/2000 7200/800
pol 15000 48 2 10800/1200/3000 10800/1200
MiniBooNE 130064 50 2 93645/10406/26013 93645/10406
credit card clients 30000 23 2 21600/2400/6000 21600/2400
MNIST 70000 28x28 10 48610/9723/11667 48610/9723
CIFAR10 60000 32x32x3 10 41665/8334/10001 41665/8334
CIFAR100 60000 32x32x3 100 41665/8334/10001 41665/8334
NCI1 4110 37 2 3328/370/412 3328/370
REDDIT-B 2000 1 2 1620/180/200 1620/180
Multi30k 31000 128x50257 Multilabel 29046/968/1000 29046/968
Note that for Multi30k, the maximum sequence length is 128 and the number of possible tokens is 50527. The classification loss is computed by comparing the predicted token against the expected one, ignoring any token beyond the length of the actual target. For a detailed description of the graph datasets, the reader can refer to Errica et al. (2020).
Appendix G Hyper-parameters for fixed and AWN models
For each data domain, we discuss the set of hyper-parameters tried during model selection for the fixed baseline and for AWN. We try our best to keep architectural choices identical with the exception of the width of the MLPs used by each model. In all experiments, we either run patience-based early stopping (Prechelt, 1998) or simply select the epoch with best validation score.
@@ G.1 Synthetic Tabular Datasets and Permuted MNIST
We report the hyper-parameter configurations tried for the fixed MLP/RNN and AWN in Table 4. In particular, the starting width of AWN under the quantile function evaluated at $k=0.9$ is approximately $256$ hidden neurons. We also noticed that a learning rate value of $0.1$ makes learning the width unstable, so we did not use it in our experiments. The total number of configurations for the fixed models is 180, for AWN is 18. We also did not impose any prior on an expected size to see if AWN recovers a similar width compared to the best fixed model (please refer to Table 1) starting from the largest configuration among the fixed models.
Table 4: Hyper-parameter configurations for standard MLP/RNN and AWN versions on tabular datasets.
Hyper-Parameter DoubleMoon Spiral SpiralHard/pol/MiniBooNE/credit c.
Batch Size 32 $[32,128]$ (MLP/RNN), 128 (AWN) $[32,128]$ (MLP/RNN), 128 (AWN)
@@
Optimizer Adam, learning rate $\in[0.1,0.01]$ (MLP/RNN), $0.01$ (AWN)
AWN Specific
@@ G.2 Image Classification Datasets
Table 5: Hyper-parameter configurations for standard MLP and AWN versions on image classification datasets.
@@
In this case, we test 5 different configurations of the fixed baseline, whereas we do not have to perform any model selection for AWN. The rate of the exponential distribution has been chosen to have a starting width of approximately $128$ neurons, which is the median value among the ones tried for the fixed model.
@@ G.3 Text Translation Tasks
Due to the cost of training a Transformer architecture from scratch on a medium size dataset like Multi30k, we use most hyper-parameters’ configurations of the base Transformer in Vaswani et al. (2017), with the difference that AWN is applied to the MLPs in each encoder and decoder layer. We train an architecture with 6 encoder and 6 decoder layers for 500 epochs, with a patience of 250 epochs applied to the validation loss. We use an Adam optimizer with learning rate 0.01, weight decay 5e-4, and epsilon 5e-9. We also introduce a scheduler that reduces the learning rate by a factor 0.9 when a plateau in the loss is reached. The number of attention heads is set to 8 and the dropout to 0.1. We try different widths for the MLPs, namely [128,256,512,1024,2048]. The embedding size is fixed to 512 and the batch size is 128. The AWN version starts with an exponential distribution rate of 0.004, corresponding to approximately 512 neurons, and a $\sigma^{\alpha}_{\ell}$=1.
@@ G.4 Graph Classification Tasks
We train the AWN version of Graph Isomorphism Network, following the setup of Errica et al. (2020) and reusing the published results. We test 12 configurations as shown in the table below. Please note that an extra MLP with 1 hidden layer is always placed before the sequence of graph convolutional layers.
Table 6: Hyper-parameter configurations for AWN versions on graph classification datasets.
@@ Appendix H Re-training a fixed network with the learned width
While on DoubleMoon AWN learns a similar number of neurons chosen by the model selection procedure and the convergence to the perfect solution over 10 runs seem much more stable, this is not the case on Spiral, SpiralHard and REDDIT-B, where AWN learns a higher total width over layers. To investigate whether the width learned by AWN on these datasets is indeed the reason behind the better performances (either a more stable training or better accuracy), we run again the experiments on the baseline networks, this time by fixing the width of each layer to the average width over hidden layers learned by AWN and reported in Table 1. We call this experiment “Fixed+”.
@@
The results suggest that by using more neurons than those selected by grid search – possibly because the validation results were identical between different configurations – it is actually possible to improve both accuracy and stability of results on Spiral and REDDIT-B, while we had a convergence issue on SpiralHard. Notice that the number of neurons learned by AWN on REDDIT-B was well outside of the range investigated by Xu et al. (2019) and Errica et al. (2020), demonstrating that the practitioner’s bias can play an important role on the final performance.
Because we are not aware of the best width achieved by the grid search approach of Errica et al. (2020) on REDDIT-B, we ran an additional experiment on fixed models, testing smaller and larger widths compared to the one selected by AWN (average of $\sim$793). What we discover is the following.
Table 7: We report REDDIT-B validation performances, averaged across the 10 outer folds, for a fixed-width DGN. The other hyper-parameters match those selected by AWN during model selection, as per this specific experiment.
# Total Width Mean Std
50 89.28 3.50
250 92.89 1.45
500 92.67 2.03
793 92.83 3.55
1000 93.06 1.60
1250 92.78 2.33
1500 91.78 2.41
Though these results are statistically comparable, the average performance obtained using AWN’s learned width seems to be close to a local optimum for the fixed-width model.
@@ Appendix I Further qualitative results on image classification datasets
Because it is well known that MNIST, CIFAR10, and CIFAR100 are datasets of increasing difficulty, we attach a figure depicting the learned number of neurons and the convergence of AWN on these datasets, following the practices highlighted in He et al. (2016) and described in Table 5 to bring the ResNet20 to convergence. Akin to the tabular datasets, we see a clear pattern where the learned number of neurons, averaged over the 10 final training runs, increases together with the task difficulty. The AWN version of the ResNet20 converges in the same amount of epochs as the fixed one.
@@
Appendix J Comparison with Neural Architecture Search (NAS)
Although complementary to our approach, we thought instructive to run a some basic NAS methods on the image classification datasets. Most NAS methods fix an upper bound on the maximum width, whereas AWN’s methodology does not. The truncation ability of AWN is not intended as a substitute for pruning but rather as a confirmation of our claims. In fact, NAS and pruning can actually be combined with AWN and do not necessarily be seen as “competing” strategies.
The results are shown in Table 8. That said, we compared the performance of AWN against simple NAS methods on the vision datasets. We tested grid search, random search (Following Wu et al. (2020)), local search (White et al., 2021), and Bayesian Optimization (Barber, 2012), fixing the same budget of 5 for all methods and width values between 0 (Linear classifier) and 512 neurons. Results are comparable to those of AWN.
Table 8: Comparison between different NAS approaches and AWN. We remind the reader that NAS is an orthogonal and complementary research direction to AWN.
Grid Search Random Search Local Search Bayes. Optim. AWN
MNIST 99.6(0.1) 99.6(0.0) 99.5 99.4 99.7(0.0)
CIFAR10 91.4(0.2) 90.1(0.5) 90.6 91.2 91.4(0.2)
CIFAR100 66.5(0.4) 64.9(1.1) 64.9 65.9 63.1(4.0)
Appendix K Non-linearity ablation study
@@
Appendix L Learned width for each layer on Graph Datasets
@@
Appendix M Quantile Ablation Study
Figure 12: We report learned width and validation accuracy trends averaged over 1-layer AWN configurations on SpiralHard. We see how a higher quantile $k$ (i.e., a better ELBO approximation) grants more stable performances. All architectures keep adapting to the task.
Appendix N Regularization Ablation Study
In Section 3, we argued there are two ways we can maximize the importance rescaling effect, namely by using bounded activation functions and by regularizing the weights. In this section, we show with an ablation that these strategies are generally not required, in the sense that AWN is still able to solve the tasks, although it might introduce a more neurons than needed. The table below compares our original results on tabular data (Table 1) with those where we did not consider ReLU6 and where we imposed an almost null regularizer on the weights ($\sigma_{\ell}^{\theta}=100$).
Table 9: We compare performances and learned width of AWN, where we removed regularization and bounded activations from hyper-parameters, with the results in the main paper.
AWN AWN - NoReg Width (AWN) Width (AWN - NoReg)
Mean (Std) Mean (Std) Mean (Std) Mean (Std)
DoubleMoon 100.0 (0.0) 100.0 (0.2) 8 (2.8) 17 (16.6)
Spiral 99.8 (0.1) 100.0 (0.0) 66 (8.7) 108 (26.8)
SpiralHard 100.0 (0.0) 100.0 (0.0) 227 (32.4) 539 (809.8)
pol 99.2 (0.1) 99.2 (0.1) 84 (11.0) 2335 (552.8)
MiniBooNE 93.2 (0.1) 93.8 (0.2) 53 (11.1) 4907 (1141)
credit card 81.8 (0.1) 81.8 (0.1) 51 (12.0) 53 (13)
We see that in all cases, not using regularization nor bounded activations leads to comparable performances but definitely higher widths. For some of the configurations tried, the width tends to grow a lot, which is why we set a maximum width of 5000 in the interest of time, although performances do not seem to decrease in spite of that. Further investigation into the results revealed that the weight regularization is what helps the most in controlling a stable behavior of the width, while the bounded activation did not have a clear impact on that. We therefore recommend some degree of regularization to make sure that the effect of the importance rescaling is properly taken into account by the optimization process.
Appendix O Wall-Clock Time Comparison
To further highlight the advantages brought by AWN, we report the average runtime (in seconds) required to complete a single model selection configuration for AWN and for the fixed model. In cases where model selection is not performed (e.g., AWN on vision tasks), we instead report the average runtime of a final run. Results for NCI1 and REDDIT are excluded from this comparison, since they were taken directly from the literature. Based on these runtimes, we compute two measures: i) the ratio of average single-run times (Single-Run Ratio), and ii) the ratio of total time required by AWN and the fixed models to complete the entire model selection process (Model Selection Ratio).
Table 10: We report average wall-clock times for AWN and the fixed model, measured either over model selection configurations or final runs when no selection is performed. We also provide two ratios: i) single-run wall-clock times and ii) total model selection time (fixed models always test at least five widths).
Dataset Avg Fixed (s) Avg AWN (s) Single-Run Ratio ($\downarrow$) Model Selection Ratio ($\downarrow$)
[…]
The results confirm that AWN’s overhead of adapting the network vanishes for larger datasets with higher feature dimensionality. Interestingly, on some of these datasets the average single-run wall time is even shorter than the fixed models. By being able to reduce hyper-parameter tuning of the width while preserving most performances, the total model selection time ratios are consistently in favor of AWN, and in principle scale favorably with the number of widths tested. Note that memory consumption scales with the width of the layers, but our methodology does not introduce any noticeable extra memory requirements.
Appendix P Impact of Removing Most Important Neurons
We complement the truncation analysis of Figure 5 by showing what happens when we remove individual neurons or a group of them, starting from the most important ones. This analysis provides further insights about how neurons behave in AWN.
Figure 13: We show the impact of removing the most important neurons, either in batch (top) or individually (bottom) on Spiral performances, using the same experiment as Figure 5.
We observe that performance quickly drops, as expected, and we reach random performance after removing the first $\sim$40 neurons out of 83. The second analysis prunes individual neurons to understand how much their absence influences performance – which can have a different effect compared to removal of a set of neurons – to see if there exist two neighboring neurons that contribute similarly to the performance drop or do not contribute at all. Removing most of the first 25 neurons out of 83 contributes majorly to performance degradation, but interestingly the removal of some neurons does not affect performances. Instead, removing any subsequent neuron has little impact (but for one) on performances, as one would expect. This suggests that we may be able to compress the neural network even more: we leave this investigation to future work.